if each edge of a cube is increased by 30% , by what percent does the volume increase?
Answer by richwmiller(17219)

v=L*W*H
1.3^3*L*W*H=1.30L*1.30W*1.30W
1.3^2=1.69
1.69*1.3=2.197 times the original
1/2.197*100% increase
old/new= percent increase